OpenStudy (goformit100):

An architect needs to determine the slope between two points on a ski lift. The two points have been identified as (15, 45) and (150, 60), where x is the horizontal distance and y is the vertical distance from the bottom of the lift. Assuming the lift runs in a straight line, what is the slope of the line between the two points? (Write your answer in simplest form, using / for a fraction bar if needed.)

OpenStudy (goformit100):

@Jonask

OpenStudy (whpalmer4):

Formula for the slope of a line between two points \((x_1,y_1)\) and \((x_2,y_2)\) is \[m = \frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

\[y_2=60,y_1=45,x_2=15,x_1=150\]

OpenStudy (anonymous):

you need to subtract the y values and divide them with the x values

OpenStudy (anonymous):

\[m=(45-60)/(150-15)=-15/135

OpenStudy (anonymous):

\[m=(45-60)/(150-15)=-15/135\]

OpenStudy (whpalmer4):

Simplest form would include reducing that fraction...